g(x)=?0 to x of cos^4tdt
then what is the value of g(x+p)
a)g(x)+g(p)
b)g(x)-G(p)

Grade:12

1 Answers

AKASH GOYAL AskiitiansExpert-IITD
13 years ago

Dear Nikhil

Your question is incomplete. But I am assuming that p is period of the function here

g(x+p)=0 to x+p∫cos4tdt= 0 to x∫cos4tdt + x to x+p ∫cos4tdt

=g(x) + 0 to p ∫cos4tdt,   [using property of periods]

=g(x) + g(p)

option a

 

All the best.                                                           

AKASH GOYAL

AskiitiansExpert-IIT Delhi

 

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively :Click here to download the toolbar..